OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (compassionate):

Hello, 2x + 3y = 5 We know the line is parallel, so it has the same slope. And we want the points 9, -3 Put the equation in y = mx + b 2x + 3y = 5 3x = -2x + 5 x = -2/3x + 5 So we know our slope is -2/3x Now plug that into the equation y = y1 = m(x = x1)